OpenStudy (idealist10):

Find the general solution of (1+x)y'+2y=sin(x)/(1+x).

OpenStudy (idealist10):

I divided (1+x) and got y'+(2/(1+x))y=sin x

OpenStudy (anonymous):

gimme a sec so i can solve it

myininaya (myininaya):

I think you should have y'+(2/(1+x))y=sin(x)/(1+x)^2

myininaya (myininaya):

if you had had (1+x)y'+2y=sin(x)*(1+x) then you could say what you said

OpenStudy (idealist10):

I see. HAHA, why am I so dumb?

OpenStudy (idealist10):

No wonder the problem won't work. hahahahaha

OpenStudy (idealist10):

Let me work it out by myself.

OpenStudy (idealist10):

I got it! Thanks a lot!

myininaya (myininaya):

np